Two friends live 25 miles apart. They walk from their homes meeting 2 1/2 hours later. One walked 2 miles per hour faster than the other. What were their two rates?
Answer by richwmiller(17219)

r*t=d
r*2.5+(r+2)*(2.5)=25
r=4 mph
r+2=6 mph
check
4*2.5+6*2.5=25
10+15=25
ok
NB 6 mph is not walking. That is jogging.
That is a mile in 10 minutes.